Andy leaves on a cross country car trip at 7:00 am. He averages 50 miles per hour. Celine plans to drive the same route, but does not leave until 8:00 am. She averages 60 miles per hour. At what time will Celine catch up with Andy?
Answer by checkley71(8403)

50T=60(T-1) OR 50T=60T-60 OR 10T=60 OR T=6 HOURS
PROOF 50*6=60(6-1) OR 300=60*5 OR 300=300